sor ma 6mH 3 Com H t 2.23MF sor vo (RUS I No initial stored energy, A Determine the state space model of this Corcuit
Answer #1

- Given I a 502 Guido 7 bulb 3 10MH V223 COMH sor vo Vc 7 22345 loop, KET 2234 ① As there are 3 energy storage elements, therex (WEEB) x (WEEi- ón (w 80:0) = ? SI WEE-B – I(WEE-B) — 1(W80.2) = ID (1os-Ios + + -] 19 = 1 frog - ljostain a lot on-Tosat node a - YtVc=vo (60w) dit +Vc = +501, - 5012 [oorn eq® = (5M) I, -(5M)12-01M) Vc +6.05M) Vo at x2 = (51) X1 -(5M)X2-(0.